OpenStudy (credmond):

Last one guys, Sorry! Can someone show me how to do this step by step!? Use implicit differentiation to find dy/dx cos(xy)=x

OpenStudy (perl):

use chain rule

OpenStudy (perl):

-sin(x*y)* ( 1* y + x * y ' ) = 1

OpenStudy (anonymous):

\[\frac{d}{dx}\cos(w) = \frac{d}{dx}(\cos(w)) \times \frac{d}{dx}(w) = -\sin(w) \cdot \frac{d}{dx}(w)\]

OpenStudy (anonymous):

First we take derivative of function, which is cos here and then we take derivative of its angle.. :)

OpenStudy (perl):

you have a mistake

OpenStudy (perl):

d/dx cos (w) = d/dw (cos(w)) * d/dx (w)

OpenStudy (perl):

d/dw (cos(w)) = - sin(w)

OpenStudy (anonymous):

I am getting you.. :) thanks for informing that.. :)

ganeshie8 (ganeshie8):

Oh fine point.

OpenStudy (credmond):

So do we have to simplify your answer perl?

OpenStudy (perl):

we can distribute

OpenStudy (credmond):

or is that the final product?

OpenStudy (perl):

since you want to solve for y ' , which is dy/dx

OpenStudy (credmond):

right.

OpenStudy (anonymous):

I think he has missed \(y\) there ? Right?

OpenStudy (anonymous):

\(-\sin(xy)[1 \cdot y + x \cdot y'] = 1\) \(-\color{red}{y} \sin(xy) - xy' \sin(xy) = 1 \)

OpenStudy (perl):

yes, woops

OpenStudy (perl):

this 'live preview' is forcing me to see double, which is distracting

OpenStudy (anonymous):

take \(-y \sin(xy)\) to other side : \[-x \cdot y' \sin(xy) = 1 + y \sin(xy) \implies y' = ??\]

OpenStudy (perl):

cos(x*y) = 1 implicit derivative: -sin(x*y)* ( 1* y + x * y ' ) = 1 - sin(xy) * y - sin(xy) * x * y ' = 1 - sin(xy) *x * y ' = 1 + y * sin(xy) y ' = ( 1 + y * sin (xy ) ) / (- x* sin xy)

OpenStudy (credmond):

Oh no worries! Okay awesome! Thank you all sooo much for all your help! You're lifesavers!! :) I wish I could give you all a million metals lol Have a great night!